OpenStudy (waheguru):

This is the equation of an arch h=-0.2d^2+3.2 How wide is the arch at the base

OpenStudy (waheguru):

How Do I solve this?

OpenStudy (waheguru):

I set h = 0 but then I was getting 4 but actually the answer says 8

OpenStudy (mathstudent55):

You are on the right track. Solve the equation with h = 0. You get a quadratic with two solutions, \(x_1\) and \(x_2\). The answer is the distance between those two solutions: \( | x_1 - x_2 | \)

OpenStudy (waheguru):

Do you mind showing?

OpenStudy (waheguru):

When you square root something, it is both positive and negative so basically I subtract 4--4 giving me 8 right? @mathstudent55

OpenStudy (waheguru):

Square Root of 16 is 4 and - 4 and by subtracting them you get 8 right?

OpenStudy (mathstudent55):

\(h=-0.2d^2+3.2 \) \(0=-0.2d^2+3.2 \) \( 0.2d^2 = 3.2 \) \(d^2 = 16\) d = 4 or d = -4 |4 - 4| = |8| = 8 or |-4 - 4| = |-8| = 8

OpenStudy (waheguru):

Thank you so much

OpenStudy (mathstudent55):

welcm